In ten years time Tom will be twice as old as Sally. Eight years ago Tom was 5 times older than Sally.Let Tom’s current age in years be x and Sally’s current age in years by y.

a. Write down and simplify two different equations connecting their ages.
b. Solve these equations simultaneously to find Tom and Sally’s current ages

Answer by mananth(16946)

tom = x
Sally =y
...
10 years later
tom = x+10
sally = y+10
...
x+10 = 2(y+10)
x+10 =2y+20
x-2y=10....................1
8 years ago
tom's age = x-8
sally's age = y-8
...
x-8=5(y-8)
x-8 = 5y-40
x-5y-40+8
x-5y=-32....................2
multiply equation by -1 and add to 2
-x+2y+x-5y=-10-32
-7y=-42
/7
y=-42/-7
y= 6 years = Sally's age now
...
...
plug value of y in equation 1
x-2*6=10
x-12=10
x= 22 tom's age now
..